# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ 4.3 Watermark verification
SEAL GCN GraphSAGE
C.ele [54] $48.90$ $50.65$ $39.35$
USAir [55] $10.56$ $49.69$ $40.07$
NS  [56] $5.06$ $64.82$ $41.69$
Yeast [52] $60.80$ $42.35$ $66.45$
Power [54] $40.55$ $52.29$ $53.04$
arXiv [57] $12.27$ $10.00$ $28.96$
PPI [58] $35.80$ $32.77$ $40.74$
Note: A watermark threshold for a dataset-architecture pair that may appear “low” (e.g., arXiv-GCN, NS-SEAL) can still be used for the purpose of ownership demonstration, because the corresponding $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ is also “low” as empirically observed in TABLES B.1, B.2, B.3.
4.4 Ownership demonstration
We now outline the process for $\mathcal{O}$ to demonstrate her ownership over $\mathcal{A}$’s model (i.e., $\mathcal{M}_{adv}$). The process uses $\mathcal{J}$ briefly outlined in §3. The process is initiated when $\mathcal{O}$ inspects $\mathcal{M}_{adv}$’s prediction API, verifies her watermark using $\mathcal{D}_{wm}$, and suspects $\mathcal{M}_{adv}$ to be plagiarised from $\mathcal{M}_{wm}$. The following steps are then initiated:
  1. 1.
    $\mathcal{O}$ accuses $\mathcal{A}$ of plagiarising her model $\mathcal{M}_{wm}$.
  2. 2.
    $\mathcal{A}$ sends $\mathcal{M}_{adv}$ to $\mathcal{J}$ for an evaluation.
  3. 3.
    $\mathcal{O}$ sends $\mathcal{D}_{wm}$, the watermark threshold corresponding to it, the code for model evaluation on $\mathcal{M}_{adv}$, and the hashes of all the files that are sent. Here, the hashes are sent to ensure that the files are not tampered with when sent via a secure communication channel.
  4. 4.
    $\mathcal{J}$ first runs a check on the hashes of the files sent. Next, $\mathcal{J}$ writes the watermark threshold onto a time-stamped public bulletin board, e.g., blockchain. A blockchain can provide the proof of anteriority in case $\mathcal{A}$ challenges $\mathcal{O}$ to have arbitrarily changed the watermark threshold later [^6] We assume $\mathcal{O}$ will send correct thresholds truthfully at the beginning.. At this stage, $\mathcal{J}$ evaluates $\mathcal{M}_{adv}$ on $\mathcal{D}_{wm}$ using the model evaluation code sent by $\mathcal{O}$.
  5. 5.
    For each data point $x$ in $\mathcal{D}_{wm}$ evaluated on $\mathcal{M}_{adv}$, soft labels are returned along with its hashes. All soft labels whose hashes are verified are collected and used to calculate the AUC score.
The ownership demonstration process ends with a comparison of the calculated AUC score against the watermark threshold sent by $\mathcal{O}$, settling the dispute between $\mathcal{O}$ and $\mathcal{A}$ with a just verdict. Figure 4 sums up the overall process from watermark data generation to ownership demonstration.
Figure 4: An overview of the watermarking process from watermark data generation to ownership demonstration.
@@ 5 Evaluation
We evaluate GENIE using 7 real-world datasets on 3 model architectures. We describe our experimental setup in §5.1. We evaluate the functionality preservation requirement of GENIE in §5.2. The robustness of GENIE is tested against 3 model extraction attacks (i.e., soft, hard, and double extraction; cf. §5.3.1) and 11 state-of-the-art backdoor defenses (i.e., 1 knowledge distillation (cf. §5.3.2), 4 model fine-tuning methods (cf. §5.3.3), 2 model compression techniques (cf. §5.3.4), 4 fine-pruning (cf. §5.3.5)). We further evaluate non-ownership piracy and efficiency requirements of GENIE in §5.4 and §5.5, respectively.
TABLE IV: Watermark verification performance (average of 10 runs) of GENIE across 3 model architectures and 7 datasets.
Dataset SEAL GCN GraphSAGE
C.ele 87.84 $\pm$ 0.46 87.60 $\pm$ 0.10 84.28 $\pm$ 0.93 88.97 $\pm$ 0.44 87.93 $\pm$ 0.43 100 $\pm$ 0.00 86.76 $\pm$ 0.68 85.71 $\pm$ 0.87 100 $\pm$ 0.00
USAir 93.19 $\pm$ 0.25 93.64 $\pm$ 0.17 92.29 $\pm$ 0.58 90.02 $\pm$ 0.52 89.35 $\pm$ 0.72 100 $\pm$ 0.00 92.44 $\pm$ 0.35 92.29 $\pm$ 0.65 100 $\pm$ 0.00
NS 98.10 $\pm$ 0.15 98.11 $\pm$ 0.23 98.70 $\pm$ 0.03 95.44 $\pm$ 0.74 96.26 $\pm$ 0.88 99.78 $\pm$ 0.00 90.90 $\pm$ 0.63 93.66 $\pm$ 0.47 99.78 $\pm$ 0.00
Yeast 97.07 $\pm$ 0.21 97.38 $\pm$ 0.16 97.69 $\pm$ 0.33 93.64 $\pm$ 0.40 91.73 $\pm$ 0.39 100 $\pm$ 0.00 89.12 $\pm$ 0.43 90.70 $\pm$ 0.43 100 $\pm$ 0.00
Power 84.41 $\pm$ 0.44 83.91 $\pm$ 0.25 88.28 $\pm$ 0.03 99.36 $\pm$ 0.17 99.12 $\pm$ 0.19 99.00 $\pm$ 0.00 87.54 $\pm$ 1.02 92.68 $\pm$ 1.06 99.00 $\pm$ 0.00
arXiv 98.14 $\pm$ 0.14 97.17 $\pm$ 0.49 98.15 $\pm$ 0.16 99.31 $\pm$ 0.04 98.78 $\pm$ 0.15 $9$$\pm$ 0.00 99.62 $\pm$ 0.01 99.32 $\pm$ 0.13 $9$$\pm$ 0.00
PPI 89.63 $\pm$ 0.12 89.45 $\pm$ 0.16 84.28 $\pm$ 1.38 95.08 $\pm$ 0.04 94.82 $\pm$ 0.05 100 $\pm$ 0.00 94.03 $\pm$ 0.09 94.31 $\pm$ 0.16 100 $\pm$ 0.00
@@ 5.1 Experimental setup
Datasets: Following prior works [25, 24], we use 7 publicly available real-world graph datasets of varying sizes and sparsities in our experiments. USAir [55] is a network of US Airlines with 332 nodes and 2,126 edges. NS [56] is a collaboration network of researchers in network science with 1,589 nodes and 2,742 edges. Yeast [52] is a protein-protein interaction network in yeast with 2,375 nodes and 11,693 edges. C.ele [54] is a neural network of C.elegans with 297 nodes and 2,148 edges. Power [54] is an electrical grid network of the western US with 4,941 nodes and 6,594 edges. arXiv [57] is a collaboration network of arXiv Astro Physics with 18,772 nodes and 198,110 edges from the popular Stanford SNAP dataset library. PPI [58] is a protein-protein interaction network with 3,890 nodes and 76,584 edges from BioGRID database.
It is worth highlighting that none of these datasets, except PPI, have node attributes. Since GENIE uses node attributes for watermark embedding, we generate node features using Node2Vec [24] for all the datasets. We follow 80-10-10 train-validation-test split of all the datasets across all our experiments. We use Adam optimizer and negative log likelihood loss for model training. Please refer Appendix A for our watermarking rates.
Models: We implement GENIE for SEAL [25] (in case of subgraph-based LP), and for GCN [8] as well as GraphSAGE [26] (in case of node representation-based LP).
SEAL: We use DGCNN as the GNN engine of SEAL. We use the default setting of DGCNN, i.e., four convolutional layers (32, 32, 32, 1 channels), a SortPooling layer (with $k=0.6$), two 1-D convolution layers (with 16, 32 output channels), and a 128-neuron dense layer. We train our models for a total of 50 epochs (for both training with or without a watermark).
GCN, SAGE: We use a 3-layer GCN and GraphSAGE models with a hidden layer of dimension 256. We use a 3-layer MLP for downstream binary classification with 256 hidden layer neurons. We train our models for a total of 400 epochs (for both training with or without a watermark).
@@
We show the robustness of GENIE against various watermark-removal techniques in the following sections. $\mathcal{A}$ might fail or succeed while trying to remove watermark from $\mathcal{M}_{wm}$. We define the success and failure of GENIE as:
@@
5.2 Functionality preserving
To evaluate the functionality-preserving nature of GENIE, we compare the performance of $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ on $\mathcal{D}_{test}$. We further assess the effectiveness of GENIE by evaluating the performance of $\mathcal{M}_{wm}$ on $\mathcal{D}_{wm}$. A high AUC score on $\mathcal{D}_{wm}$ indicates that the model has successfully learned to associate the watermark patterns with the predetermined incorrect predictions; enabling reliable watermark detection and ownership verification. TABLE IV presents $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$, $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ across all 7 datasets and 3 model architectures considered in our work. The scores reported here are averaged from 10 runs using different random seeds (to mitigate potential seed influence and capture real the mean performance of GENIE). We observe that there is less than $2\%$ decrease in the AUC score on $\mathcal{D}_{test}$ due to watermarking across all model architectures for all datasets. These results allow us to claim that GENIE is functionality preserving. Since $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ is more than $80\%$ across all model architectures for all datasets, we can claim that $\mathcal{M}_{wm}$ has successfully learned $\mathcal{F}_{wm}$. Interestingly, we observe an increase in AUC score on $\mathcal{D}_{test}$ after watermarking in a few cases, it could be a result of the watermarking embedding process acting as a regularizer, which reduces the overfitting nature of the model.
5.3 Robustness
To ensure the robustness and reliability of a watermarking scheme, it is crucial to assess its resilience against potential attempts by $\mathcal{A}$ to watermark. Since $\mathcal{A}$ cannot directly verify the presence of watermark in $\mathcal{M}_{wm}$, $\mathcal{A}$ may resort to various techniques (e.g., model extraction, model pruning, fine-tuning) to eliminate or degrade the embedded watermark. We conduct a series of evaluations to extensively examine the effectiveness of GENIE against a diverse set of robustness tests and potential attacks.
Due to the page length limit, here we present the results from robustness tests on GENIE for GCN only. We detail the corresponding results for SEAL and GraphSAGE in Appendix D.

### v3

@@ 4.3.2 Dynamic Watermark Thresholding (DWT)
SEAL GCN GraphSAGE NeoGNN
C.ele [44] $48.90$ $50.65$ $39.35$ $38.42$
USAir [45] $10.56$ $49.69$ $40.07$ $18.02$
NS  [46] $5.06$ $64.82$ $41.69$ $41.44$
Yeast [41] $60.80$ $42.35$ $66.45$ $12.63$
Power [44] $40.55$ $52.29$ $53.04$ $54.85$
arXiv [47] $12.27$ $10.00$ $28.96$ $16.22$
PPI [48] $35.80$ $32.77$ $40.74$ $36.76$
4.4 Ownership Demonstration (OD)
TABLE IV: Watermark verification performance (average of 10 runs) of Genie across 4 model architectures and 7 datasets.
Dataset SEAL GCN GraphSAGE NeoGNN
$AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ (%) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%)
C.ele 87.84 $\pm$ 0.46 87.60 $\pm$ 0.10 84.28 $\pm$ 0.93 88.97 $\pm$ 0.44 87.93 $\pm$ 0.43 100 $\pm$ 0.00 86.76 $\pm$ 0.68 85.71 $\pm$ 0.87 100 $\pm$ 0.00 89.03 $\pm$ 0.71 88.94 $\pm$ 1.20 100 $\pm$ 0.00
USAir 93.19 $\pm$ 0.25 93.64 $\pm$ 0.17 92.29 $\pm$ 0.58 90.02 $\pm$ 0.52 89.35 $\pm$ 0.72 100 $\pm$ 0.00 92.44 $\pm$ 0.35 92.29 $\pm$ 0.65 100 $\pm$ 0.00 95.81 $\pm$ 0.81 94.57 $\pm$ 1.45 100 $\pm$ 0.00
NS 98.10 $\pm$ 0.15 98.11 $\pm$ 0.23 98.70 $\pm$ 0.03 95.44 $\pm$ 0.74 96.26 $\pm$ 0.88 99.78 $\pm$ 0.00 90.90 $\pm$ 0.63 93.66 $\pm$ 0.47 99.78 $\pm$ 0.00 99.93 $\pm$ 0.02 99.80 $\pm$ 0.14 100 $\pm$ 0.00
Yeast 97.07 $\pm$ 0.21 97.38 $\pm$ 0.16 97.69 $\pm$ 0.33 93.64 $\pm$ 0.40 91.73 $\pm$ 0.39 100 $\pm$ 0.00 89.12 $\pm$ 0.43 90.70 $\pm$ 0.43 100 $\pm$ 0.00 97.78 $\pm$ 0.57 97.54 $\pm$ 0.19 100 $\pm$ 0.00
Power 84.41 $\pm$ 0.44 83.91 $\pm$ 0.25 88.28 $\pm$ 0.03 99.36 $\pm$ 0.17 99.12 $\pm$ 0.19 99.00 $\pm$ 0.00 87.54 $\pm$ 1.02 92.68 $\pm$ 1.06 99.00 $\pm$ 0.00 99.96 $\pm$ 0.02 99.94 $\pm$ 0.04 100 $\pm$ 0.00
arXiv 98.14 $\pm$ 0.14 97.17 $\pm$ 0.49 98.15 $\pm$ 0.16 99.31 $\pm$ 0.04 98.78 $\pm$ 0.15 $9$$\pm$ 0.00 99.62 $\pm$ 0.01 99.32 $\pm$ 0.13 $9$$\pm$ 0.00 99.92 $\pm$ 0.01 99.91 $\pm$ 0.01 94.22 $\pm$ 3.99
PPI 89.63 $\pm$ 0.12 89.45 $\pm$ 0.16 84.28 $\pm$ 1.38 95.08 $\pm$ 0.04 94.82 $\pm$ 0.05 100 $\pm$ 0.00 94.03 $\pm$ 0.09 94.31 $\pm$ 0.16 100 $\pm$ 0.00 97.43 $\pm$ 0.16 97.44 $\pm$ 0.11 97.64 $\pm$ 1.77
We now outline the process for $\mathcal{O}$ to demonstrate her ownership over $\mathcal{A}$’s model (i.e., $\mathcal{M}_{adv}$). OD uses $\mathcal{J}$ briefly outlined in §3. Genie has a two-step OD procedure that involves a model registration step and a dispute resolution step.
Model registration: As a preemptive step of OD, $\mathcal{O}$ first sends $\mathcal{G}$ to $\mathcal{J}$ to procure $\mathcal{D}_{wm}$, addressing the problem of malicious plaintiff outlined in [8]. $\mathcal{J}$ then writes the cryptographic hash of $\mathcal{D}_{wm}$ onto a time-stamped public bulletin board (e.g., blockchain) to provide the proof of anteriority in case of a dispute. We call this preemptive step model registration since it is analogous to patent registration common in the protection of IP rights [49, 50]. The procured $\mathcal{D}_{wm}$ will then be used for embedding the watermark into $\mathcal{M}$, i.e., train an untrained $\mathcal{M}$ to be $\mathcal{M}_{wm}$ by $\mathcal{O}$.
Dispute resolution: When a dispute arises, OD involves the following steps: (1) $\mathcal{O}$ accuses $\mathcal{A}$ of plagiarising her model $\mathcal{M}_{wm}$; (2) $\mathcal{A}$ sends $\mathcal{M}_{adv}$ to $\mathcal{J}$ for an evaluation; (3) $\mathcal{O}$ sends $\mathcal{D}_{wm}$ and the hashes of all the files. Here, the hashes are sent via a secure communication channel to ensure that the files are not tampered with; (4) $\mathcal{J}$ runs a check on the hashes of the files sent. Next, $\mathcal{J}$ checks the record of $\mathcal{D}_{wm}$ in the public bulletin board. If a matching record is found, $\mathcal{J}$ first calculates the watermark threshold $t$, and evaluates $\mathcal{M}_{adv}$ on $\mathcal{D}_{wm}$ to get $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$. The OD ends with a comparison of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ against $t$, settling the dispute between $\mathcal{O}$ and $\mathcal{A}$ with a just verdict. On the other hand, if a record is not found, the dispute resolves in $\mathcal{O}$’s defeat.
@@ 5 Evaluation
We evaluate Genie using 7 real-world datasets on 3 model architectures. We describe our experimental setup in §5.1. We evaluate the functionality preservation requirement of Genie in §5.2 and compare our watermark embedding method with previous methods in §5.3. The robustness of Genie is tested against 3 model extraction attacks (i.e., soft, hard, and double extraction; cf. §5.4.1) and 11 state-of-the-art backdoor defenses (i.e., 1 knowledge distillation (cf. §5.4.2), 4 model fine-tuning methods (cf. §5.4.3), 2 model compression techniques (cf. §5.4.4), 4 fine-pruning (cf. §5.4.5)). We further evaluate non-ownership piracy and efficiency requirements of Genie in §5.5 and §5.7, respectively. Finally, we discuss a defense strategy to counter adaptive attackers in §5.6.
@@ 5.1 Experimental setup
Datasets: Following prior works [19, 52], we use 7 publicly available real-world graph datasets of varying sizes and sparsities in our experiments: USAir, NS, Yeast, C.ele, Power, arXiv and PPI (cf. Appendix A.1 for dataset details).
It is worth highlighting that none of these datasets, except PPI, have node attributes. However, GNN training requires node features (using Node2vec [52] or random initialization at minimum). We used Node2vec to generate node features for all datasets for optimal $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$. We follow an 80-10-10 train-validation-test split of all the datasets across all our experiments. We use Adam optimizer and negative log likelihood loss for model training. Please refer Appendix A for our watermarking rates.
Models: We implement Genie for SEAL [19] (SotA subgraph-based LP method) and for NeoGNN [53] (SotA node-representation based LP method). We also implement Genie for GCN [20] and GraphSAGE [54] that are widely used GNN architectures (cf. Appendix A.2 for further details).
@@
5.2 Functionality preserving
To evaluate Genie’s functionality-preserving nature, we compare the performance of $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ on $\mathcal{D}_{test}$. We assess Genie’s effectiveness by evaluating $\mathcal{M}_{wm}$’s performance on $\mathcal{D}_{wm}$. We establish a strict threshold of 2% drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ as the criterion for a watermarking scheme to be considered functionality-preserving. A high AUC score on $\mathcal{D}_{wm}$ indicates that the model has successfully learned to associate watermark patterns with predetermined incorrect predictions, enabling reliable watermark detection and ownership verification. We set 80% as the minimum $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ for reliable watermark detection. TABLE IV presents $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$, $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$, and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ across all 7 datasets and 4 model architectures considered in our work. The reported scores are averaged from 10 runs using different random seeds to mitigate potential seed influence and capture Genie’s real mean performance. Our results show less than a 2% decrease in AUC scores on $\mathcal{D}_{test}$ due to watermarking across all model architectures and datasets, allowing us to claim that Genie is functionality preserving. Furthermore, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ exceeds 80% across all model architectures and datasets, indicating that $\mathcal{M}_{wm}$ has successfully learned $\mathcal{F}_{wm}$. Interestingly, we observe an increase in AUC scores on $\mathcal{D}_{test}$ after watermarking in some cases. This improvement could result from the watermark embedding process acting as a regularizer, potentially reducing the model’s overfitting tendency.
5.3 Comparison among embedding methods
We compare our watermark embedding method with 4 baselines: (1) Fine-Tuning: Xu et al. [12] fine-tunes a clean pre-trained model on $\mathcal{D}_{wm}$. We fine-tune for 50 epochs to avoid over-fitting on $\mathcal{D}_{wm}$. (2) Data Poisoning: Adi et al. [5] adds trigger samples to the training set to embed the watermark. (3) Uniform Loss: Minimizing the sum of loss functions ($\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$). (4) MGDA [56]: It computes scaling coefficients $\alpha_{1},\alpha_{2}$ for losses $\mathcal{L}_{1},\mathcal{L}_{2}$, respectively by minimizing the weighted sum of losses ($\mathcal{L}=\alpha_{1}\mathcal{L}_{1}+\alpha_{2}\mathcal{L}_{2}~{}|~{}\alpha_{1}+\alpha_{2}=1$). MGDA also provably converges [^5] The convergence criterion holds when all training parameters lie within the model, which is the case for Genie. Consequently, optimization methods considered in works such as [57] are inapplicable. to a solution from Pareto-optimal set [58], making it ideal for benchmarking Genie’s watermark embedding method.
We define that a practical watermark embedding method ensures a less than 2% drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ and achieves a minimum of 80% $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (cf. §5.2). Our empirical results in TABLE V reveal that the method by Xu et al. [12] violates the functionality-preserving property across all datasets, with a drop of more than 2% in model utility. While the method by Adi et al. [5] maintains functionality preservation, it performs poorly on $\mathcal{D}_{wm}$ for 4 out of 7 datasets. The uniform loss approach violates the functionality-preserving property for 3 out of 7 datasets. Despite MGDA [56] performing on par with or better than Genie in some cases, it fails to preserve functionality for large datasets (i.e., arxiv, PPI). This comparative analysis demonstrates the superiority of Genie’s watermark embedding method over existing approaches. Genie consistently preserves functionality while achieving high performance on $\mathcal{D}_{wm}$ across all datasets.
TABLE V: Comparison among watermark embedding methods.
Dataset C.ele USAir NS Yeast Power arXiv PPI
No Watermark $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ (%) 87.90 89.62 96.00 93.45 99.54 99.28 95.83
Xu et al. [12] $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) 53.41 33.32 78.57 49.80 74.74 20.88 26.28
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) 100 99.32 95.56 100 91.50 100 100
Adi et al. [5] $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) 88.50 91.45 96.23 94.06 99.28 99.06 95.95
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) 90.62 74.51 96.67 53.25 98.00 4.38 21.58
Uniform Loss ($\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{2}$) $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) 88.37 80.53 96.44 92.52 99.20 93.29 91.45
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) 100 98.83 99.78 100 99.00 100 100
MGDA [56] $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) 88.57 89.93 95.76 92.82 99.04 96.40 93.81
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) 100 100 99.78 100 99.00 100 100
Genie’s approach $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ (%) 86.93 88.34 96.59 91.46 98.92 98.13 94.67
$AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ (%) 100 100 99.77 100 99.00 100 100
Note: The highest and the second highest $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ values are bold and underlined, respectively. This table shows our ablation done using GCN model.
5.4 Robustness
@@
To ensure the robustness and reliability of a watermarking scheme, it is crucial to assess its resilience against potential attempts by $\mathcal{A}$ to watermark. Since $\mathcal{A}$ cannot directly verify the presence of watermark in $\mathcal{M}_{wm}$, $\mathcal{A}$ may resort to various techniques (e.g., model extraction, model pruning, fine-tuning) to eliminate or degrade the embedded watermark. We show the robustness of Genie against various watermark-removal techniques in the following sections. $\mathcal{A}$ might fail or succeed while trying to remove the watermark from $\mathcal{M}_{wm}$. We define the success and failure of Genie as:
@@
Due to the page length limit, here we present the results from robustness tests on Genie for GCN only. We detail the corresponding results for other models in Appendix D.